OpenStudy (anonymous):

Find ALL fifth roots of -243

Nnesha (nnesha):

\[\huge\rm \sqrt[5]{-243}\] this is your question right

OpenStudy (anonymous):

243 = 3^5

OpenStudy (anonymous):

so... \(\large \sqrt[5]{-3^5} =?\)